# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ 3.2 Upwind discretization of Godunov–Powell source term
$+\frac{1}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\frac{-\mathscr{V}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)}{\mathscr{V}_{2,j+\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)-\mathscr{V}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)}\left({\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{+}\right)-{\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-}\right)\right)\right.$
@@
$+\left.\frac{\mathscr{V}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)}{\mathscr{V}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)-\mathscr{V}_{2,j-\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)}\left({\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)-{\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{-}\right)\right)\right].$
Note that “$\text{div}_{\text{within}}\boldsymbol{B}^{n}$” means the divergence within the cell $I_{ij}$, which becomes zero because of locally divergence-free property, while “$\text{div}_{\text{faces}}\boldsymbol{B}^{n}$” involves divergence error across the cell interfaces, which ia consistent with the discretized divergence in Godunov-Powell source term.
@@ 3.3 Convex decomposition of cell averages
In the framework of provably PP high-order schemes, a key ingredient is the decomposition of the cell averages of the numerical solution into a convex combination of the solution values at certain quadrature points. The convex decomposition determines not only the theoretical PP CFL condition of the resulting scheme, but also the points to perform the PP limiter, as will be shown in the Section 3.5.
Let $\left\{x_{i}^{(\mu)}\right\}_{\mu=1}^{q}$ and $\left\{y_{j}^{(\mu)}\right\}_{\mu=1}^{q}$ denote the $q$-point Gauss-Legendre quadrature nodes with $2q-1\geq 2k$ in $\left[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}}\right]$ and $\left[y_{j-\frac{1}{2}},y_{j+\frac{1}{2}}\right]$, respectively, and $\left\{\omega_{\mu}^{G}\right\}_{\mu=1}^{q}$ be the associated weights. Denote $\left\{\hat{x}_{i}^{(\mu)}\right\}_{\mu=1}^{L}$ and $\left\{\hat{y}_{j}^{(\mu)}\right\}_{\mu=1}^{L}$ as the $L$-point Gauss-Lobatto quadrature nodes with $2L-3\geq k$ in $\left[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}}\right]$ and $\left[y_{j-\frac{1}{2}},y_{j+\frac{1}{2}}\right]$, respectively, and $\left\{\hat{\omega}_{\mu}^{GL}\right\}_{\mu=1}^{L}$ are the associated weights with $\hat{\omega}_{1}^{GL}=\hat{\omega}_{L}^{GL}=\frac{1}{L(L-1)}$. Let $a_{1}$ and $a_{2}$ denote the maximum characteristic speeds in x- and y-directions.
According to [54, 53], a general feasible convex decomposition for polynomial space $\mathbb{P}^{k}\left(I_{ij}\right)$ on Cartesian mesh can be written as
@@
and it simultaneously satisfies the following three conditions:
@@
$\mathbb{S}_{ij}^{\text{boundary}}:={\left\{\left(x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)\right\}}_{\mu=1}^{q}\bigcup{\left\{\left(x_{i-\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)\right\}}_{\mu=1}^{q}\bigcup{\left\{\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-},\right)\right\}}_{\mu=1}^{q}\bigcup{\left\{\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+},\right)\right\}}_{\mu=1}^{q},$
and the cell average of numerical solution is decomposed into a convex combination of solution values at the quadrature points belonging to $\mathbb{S}_{ij}:=\mathbb{S}_{ij}^{\text{boundary}}\bigcup\mathbb{S}_{ij}^{\text{internal}}$.
Based on the tensor product of the $L$-point Gauss-Lobatto quadrature with $L=\lceil\frac{k+3}{2}\rceil$ and the $q$-point Gauss quadrature with $q=k+1$, Zhang and Shu in [36] proposed a classic convex decomposition, which is listed as follows.
Zhang-Shu convex decomposition.
@@
Cui et al. [54, 53] discovered an optimal convex decomposition allowing much fewer nodes and larger time step-size, as shown below for the case of $\mathbb{P}^{2}$ or $\mathbb{P}^{3}$ space.
Optimal convex decomposition.
@@
$+\frac{\kappa_{2}}{2}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\boldsymbol{U}_{ij}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-},\right)+\boldsymbol{U}_{ij}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)\right]$
@@
In Figure 1, we show the nodes of optimal convex decomposition and classic convex decompositions for the polynomial space $\mathbb{P}^{2}\left(I_{ij}\right)$, on the square grid, in the case of $\frac{\Delta x}{a_{1}}=\frac{\Delta y}{a_{2}}$.
Figure 1: Boundary nodes (blue) and internal nodes (red) of the classic convex decompositions (left) and the optimal convex decomposition (right) for the polynomial space $\mathbb{P}^{2}\left(I_{ij}\right)$, on a rectangular cell, in the case of $\frac{\Delta x}{a_{1}}=\frac{\Delta y}{a_{2}}$.
3.4 Positivity-preserving limiter
Let $\boldsymbol{U}_{ij}^{n}\left(\boldsymbol{x}\right)$ denote the approximate solution polynomial in $I_{ij}$ at the $n$-th time level after the OE procedure, and $\bar{\boldsymbol{U}}_{ij}^{n}$ be the cell average of $\boldsymbol{U}_{ij}^{n}\left(\boldsymbol{x}\right)$ over $I_{ij}$. Following the analysis in [44], the OEDG method with the PP HLL flux and properly discretized Godunov-Powell source term preserves the positivity for the updated cell averages under a CFL condition, if the OEDG solution polynomials satisfy
${\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G},\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij}$ (50)
where $\mathbb{S}_{ij}$ is all the nodes of convex decomposition (47).
The OE procedure, although enhancing the numerical stability for strong shocks, does not necessarily guarantee the condition (50), which should be enforced by a simple PP limiter. For readers’ convenience, here we review the PP limiter [37, 55, 44] briefly. The PP limiting procedure mainly consists of two steps. First, modify the density to enforce the positivity by
@@ 3.4 Positivity-preserving limiter
Here $\epsilon_{1}=\min\left\{10^{-13},\bar{\rho}_{ij}^{n}\right\}$ and $\epsilon_{2}=\min\left\{10^{-13},\mathcal{E}\left(\bar{\boldsymbol{U}}_{ij}^{n}\right)\right\}$ are introduced to avoid rounding error [44]. Note that the modified polynomial solution has the same cell average $\bar{\boldsymbol{U}}_{ij}^{n}$ and holds
$\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G}_{\epsilon}=\left\{\boldsymbol{U}:\rho\geq\epsilon_{1},\mathcal{E}\left(\boldsymbol{U}\right)\geq\epsilon_{2}\right\},\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij}.$ (53)
It is worth mentioning that such a limiter does not reduce the high-order accuracy of the DG schemes [36, 37]. The PP limiter is performed right after each OE procedure.
3.5 Algorithmic flow of PP LDF OEDG scheme
Based on the LDF OEDG scheme, properly discretized Godunov-Powell source term and the convex combination of cell averages, we construct the provably PP LDF OEDG scheme as follows.
@@ 3.5 Algorithmic flow of PP LDF OEDG scheme
Initialize $\left\{\bar{\boldsymbol{U}}_{ij}^{0}\right\}$ and $\left\{\boldsymbol{U}_{ij}^{0}\left(\boldsymbol{x}\right)\right\}$ by the $L^{2}$-projection of initial data onto $\mathbb{V}_{U}^{k}$, see (LABEL:Equ:2DinitprojR)-(LABEL:Equ:2DinitprojQ). $\left\{\bar{\boldsymbol{U}}_{ij}^{0}\right\}\in\mathcal{G}$ due to the convexity of $\mathcal{G}$.
@@
Given admissible cell averages $\left\{\bar{\boldsymbol{U}}_{ij}^{n}\right\}$ and approximate polynomial solution $\boldsymbol{U}_{\sigma}^{n}\left(\boldsymbol{x}\right)\in\mathbb{V}_{U}^{k}$ defined by $\boldsymbol{U}_{ij}^{n}\left(\boldsymbol{x}\right)$, perform PP limiter to modify the polynomials $\left\{\boldsymbol{U}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ such that the modified polynomials $\left\{\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ satisfy
$\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G},\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij},$ (54)
where $\mathbb{S}_{ij}$ is all the nodes of convex decomposition (47). The PP limiting procedure can refer to Section 3.4.
@@
Let $\widetilde{\boldsymbol{U}}_{\sigma}^{n}\left(\boldsymbol{x}\right)\in\mathbb{V}_{U}^{k}$ denote the discontinuous piecewise polynomial solution defined by $\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)$. For one-order forward Euler method, update the cell averages by the scheme
$\bar{\boldsymbol{U}}_{ij}^{n+1}=\bar{\boldsymbol{U}}_{ij}^{n}+\Delta t_{n}\boldsymbol{L}_{ij}\left(\widetilde{\boldsymbol{U}}_{\sigma}^{n}\right).$ (55)
As will be shown in Theorem 1, the scheme (55) can ensure $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$.
@@
For one-order forward Euler method, evolve the high-order “moments” of the polynomials $\left\{\boldsymbol{U}_{ij}^{n+1}\left(\boldsymbol{x}\right)\right\}$ by using the LDF OEDG scheme detailed in Section 2 to build the discontinuous piecewise polynomial function $\boldsymbol{U}_{\sigma}^{n+1}\left(\boldsymbol{x}\right)\in\mathbb{V}_{U}^{k}$.

### v2

@@ 3.2 Upwind discretization of Godunov–Powell source term
$\approx\frac{1}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\bigg{[}\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\big{(}x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\big{)}\right)\omega_{\text{jump}}^{-}\left[\!\left[\left(B_{1}\right)_{\sigma}^{n}\right]\!\right]_{i+\frac{1}{2}}(y_{j}^{(\mu)})+$
$\qquad\qquad\qquad~{}~{}\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\big{(}x_{i+\frac{1}{2}}^{+},y_{j}^{(\mu)}\big{)}\right)\omega_{\text{jump}}^{+}\left[\!\left[\left(B_{1}\right)_{\sigma}^{n}\right]\!\right]_{i+\frac{1}{2}}(y_{j}^{(\mu)})\bigg{]},$
where $\left[\!\left[\left(B_{1}\right)_{\sigma}^{n}\right]\!\right]_{i+\frac{1}{2}}(y)$ is the jump across the interface $x=x_{i+\frac{1}{2}}$, and $\omega_{\text{jump}}^{-}$ (resp. $\omega_{\text{jump}}^{+}$) represents the weight assigned to the cell $I_{ij}$ (resp. $I_{i+1,j}$). These two weights satisfy $\omega_{\text{jump}}^{-}+\omega_{\text{jump}}^{+}=1$, as we have discussed in (46). In this work, we adopt the following wave-dependent weights $\omega_{\text{jump}}^{-}$ and $\omega_{\text{jump}}^{+}$:
@@
$\omega_{\text{jump}}^{-}=\frac{-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)},\qquad\omega_{\text{jump}}^{+}=\frac{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}.$
Let $\mathscr{V}_{1,l}\left(x_{i+\frac{1}{2}},y_{j}^{(\mu)}\right)$ and $\mathscr{V}_{1,r}\left(x_{i+\frac{1}{2}},y_{j}^{(\mu)}\right)$ denote the (estimated) minimum and maximum wave speeds at the interface $x=x_{i+\frac{1}{2}}$, respectively. Recall that
$\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)=\min\left\{0,\mathscr{V}_{1,l}\left(x_{i+\frac{1}{2}},y_{j}^{(\mu)}\right)\right\},\quad\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)=\max\left\{0,\mathscr{V}_{1,r}\left(x_{i+\frac{1}{2}},y_{j}^{(\mu)}\right)\right\}.$
Figure 2 shows the three HLL wave patterns at the interface $x=x_{i+\frac{1}{2}}$. The corresponding weights are listed below:
  • 1.
    For case (a), $\omega_{\text{jump}}^{-}=\frac{-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}$, $\omega_{\text{jump}}^{+}=0$;
  • 2.
    For case (b), $\omega_{\text{jump}}^{-}=\frac{-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}$, $\omega_{\text{jump}}^{+}=\frac{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}$;
  • 3.
    For case (c), $\omega_{\text{jump}}^{-}=0$, $\omega_{\text{jump}}^{+}=\frac{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}$.
Such choices reflect an “upwind” discretization of the Godunov–Powell source term. It is worth mentioning that this discretization is actually motivated by rigorous PP analysis, which will become clear in the proof of Theorem 1 in section 3.6.
Figure 2: The three HLL wave patterns at the interface $x=x_{i+\frac{1}{2}}$ with the (estimated) minimum wave speed $\mathscr{V}_{1,l}\left(x_{i+\frac{1}{2}},y_{j}^{(\mu)}\right)$ and the (estimated) maximum wave speed $\mathscr{V}_{1,r}\left(x_{i+\frac{1}{2}},y_{j}^{(\mu)}\right)$.
@@ 3.3 Convex decomposition of cell averages
In the framework [39, 57, 58] of constructing provably PP high-order schemes, a key ingredient is the decomposition of the cell averages of the numerical solution into a convex combination of the point values at certain quadrature points. The convex decomposition determines not only the theoretical PP CFL condition of the resulting scheme, but also identifies the points at which to perform the PP limiter, as will be shown in section 3.4.
Recall that $\{x_{i}^{(\mu)}\}_{\mu=1}^{q}$ and $\{y_{j}^{(\mu)}\}_{\mu=1}^{q}$ denote the Gauss quadrature nodes, and $\{\omega_{\mu}^{G}\}_{\mu=1}^{q}$ are the associated weights in the DG schemes in (9). Let $a_{1}$ and $a_{2}$ represent the (estimated) maximum wave speeds in the $x$- and $y$-directions, respectively. According to [57, 58], a general feasible convex decomposition for polynomial space $\mathbb{P}^{k}(I_{ij})$ on a rectangular cell can be expressed as
@@
which should simultaneously satisfy the following three conditions:
@@
$\mathbb{S}_{ij}^{\text{boundary}}:={\left\{\left(x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)\right\}}_{\mu=1}^{q}\bigcup{\left\{\left(x_{i-\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)\right\}}_{\mu=1}^{q}\bigcup{\left\{\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-}\right)\right\}}_{\mu=1}^{q}\bigcup{\left\{\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)\right\}}_{\mu=1}^{q}.$
Based on (50), the cell average $\bar{\boldsymbol{U}}_{ij}$ is decomposed into a convex combination of solution values at the quadrature points belonging to
$\mathbb{S}_{ij}:=\mathbb{S}_{ij}^{\text{boundary}}\bigcup\mathbb{S}_{ij}^{\text{internal}},$
at which the PP limiter will be performed; see section 3.4.
Denote $\left\{\hat{x}_{i}^{(\mu)}\right\}_{\mu=1}^{L}$ and $\left\{\hat{y}_{j}^{(\mu)}\right\}_{\mu=1}^{L}$ as the $L$-point Gauss–Lobatto quadrature nodes with $L=\lceil\frac{k+3}{2}\rceil$ in $\left[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}}\right]$ and $\left[y_{j-\frac{1}{2}},y_{j+\frac{1}{2}}\right]$, respectively, and $\left\{\hat{\omega}_{\mu}^{GL}\right\}_{\mu=1}^{L}$ are the associated weights for the interval $[-\frac{1}{2},\frac{1}{2}]$ with $\hat{\omega}_{1}^{GL}=\hat{\omega}_{L}^{GL}=\frac{1}{L(L-1)}$. Based on the tensor product of the $L$-point Gauss–Lobatto quadrature and the $q$-point Gauss quadrature, Zhang and Shu [39] proposed the following classic convex decomposition.
Zhang–Shu convex decomposition:
@@
Cui, Ding, and Wu [57, 58] discovered an optimal convex decomposition involving much fewer nodes and allowing the mildest PP time step-size, as shown below for the cases of $\mathbb{P}^{2}$ and $\mathbb{P}^{3}$ spaces. For the optimal convex decomposition for general $\mathbb{P}^{k}$ spaces with $k\geq 4$, see [58].
Optimal convex decomposition for $\mathbb{P}^{2}$ and $\mathbb{P}^{3}$:
@@
$+\frac{\kappa_{2}}{2}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\boldsymbol{U}_{ij}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-}\right)+\boldsymbol{U}_{ij}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)\right]$
@@
3.4 Positivity–preserving limiter
Let $\boldsymbol{U}_{ij}^{n}\left(\boldsymbol{x}\right)$ denote the OEDG solution in $I_{ij}$ at the $n$th time level after the LDF OE procedure, and let $\bar{\boldsymbol{U}}_{ij}^{n}$ be its cell average over $I_{ij}$. As it will be shown in section 3.6, the OEDG method with the PP HLL flux and upwind discrete Godunov–Powell source term preserves the positivity for the updated cell averages under a CFL condition, if the OEDG solution satisfies
${\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G}\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij}$ (53)
where $\mathbb{S}_{ij}$ is the set of all nodes involved in the adopted convex decomposition (50).
The LDF OE procedure, although enhancing the numerical stability for strong shocks, does not necessarily guarantee the condition (53), which should be enforced by a simple PP limiter. For readers’ convenience, here we briefly review the PP limiter [40, 42, 48]. The PP limiting procedure mainly consists of two steps. First, modify the density to enforce the positivity by
@@ 3.4 Positivity–preserving limiter
Here $\epsilon_{1}=\min\left\{10^{-13},\bar{\rho}_{ij}^{n}\right\}$ and $\epsilon_{2}=\min\left\{10^{-13},\mathcal{E}\left(\bar{\boldsymbol{U}}_{ij}^{n}\right)\right\}$ are introduced to avoid the impact of rounding errors on PP property [48]. Note that the modified solution has the same cell average $\bar{\boldsymbol{U}}_{ij}^{n}$ on $I_{ij}$ and holds
$\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G}_{\epsilon}=\left\{\boldsymbol{U}:\rho\geq\epsilon_{1},\mathcal{E}\left(\boldsymbol{U}\right)\geq\epsilon_{2}\right\}\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij}.$ (56)
It is worth mentioning that such a limiter does not reduce the high-order accuracy of the DG schemes [39, 40]. The PP limiter is performed right after each LDF OE procedure, and it maintains the LDF property for the magnetic field.
3.5 Algorithmic flow of PP LDF OEDG schemes
Based on the LDF OEDG method, the proper upwind discretization of Godunov–Powell source term, and the PP limiter, we construct the PP LDF OEDG schemes as follows. For clarity, we focus on the detailed descriptions of the schemes with the forward Euler time discretization.
@@ 3.5 Algorithmic flow of PP LDF OEDG schemes
Initialize the approximate solution $\boldsymbol{U}_{h}^{0}\left(\boldsymbol{x}\right)$ by the $L^{2}$-projection of initial data of $\boldsymbol{R}$ onto the space $\mathbb{V}_{R}^{k}$ and the $L^{2}$-projection of initial data of $\boldsymbol{Q}$ onto the LDF space $\mathbb{V}_{Q}^{k}$. The cell average $\bar{\boldsymbol{U}}_{ij}^{0}\in\mathcal{G}$, because of the convexity of $\mathcal{G}$.
@@
Given the approximate solution $\boldsymbol{U}_{h}^{n}\left(\boldsymbol{x}\right)$ with admissible cell averages $\left\{\bar{\boldsymbol{U}}_{ij}^{n}\right\}$, perform LDF OE procedure
$\boldsymbol{R}_{\sigma}^{n}\left(\boldsymbol{x}\right)=\mathcal{F}_{\text{OE}}^{R}\boldsymbol{R}_{h}^{n}\left(\boldsymbol{x}\right),\quad\boldsymbol{Q}_{\sigma}^{n}\left(\boldsymbol{x}\right)=\mathcal{F}_{\text{OE}}^{Q}\boldsymbol{Q}_{h}^{n}\left(\boldsymbol{x}\right),$ (57)
to obtain the OEDG solution $\boldsymbol{U}_{\sigma}^{n}\left(\boldsymbol{x}\right)$.
@@
Given the OEDG solution $\boldsymbol{U}_{\sigma}^{n}\left(\boldsymbol{x}\right)$ with the admissible cell averages $\left\{\bar{\boldsymbol{U}}_{ij}^{n}\right\}$, define the polynomial vector of $\boldsymbol{U}_{\sigma}^{n}\left(\boldsymbol{x}\right)$ on $I_{ij}$ as $\boldsymbol{U}_{ij}^{n}\left(\boldsymbol{x}\right)$, and perform PP limiter (section 3.4) to modify $\left\{\boldsymbol{U}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ such that the modified polynomial vector $\left\{\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ satisfies
$\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G}\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij},$ (58)
where $\mathbb{S}_{ij}$ is the set of all the nodes of the adopted convex decomposition (50). We opt to use the optimal convex decomposition (52) in our computations.
@@
Let $\widetilde{\boldsymbol{U}}_{\sigma}^{n}\left(\boldsymbol{x}\right)$ denote the piecewise polynomial solution defined by $\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)$. Perform an SSP Runge–Kutta time discretization to evolve the OEDG solution. For clarity, we only provide specific details for the forward Euler method. The cell averages $\{\bar{\boldsymbol{U}}_{ij}^{n+1}\}$ are updated by utilizing the following scheme
$\bar{\boldsymbol{U}}_{ij}^{n+1}=\bar{\boldsymbol{U}}_{ij}^{n}+\Delta t_{n}\boldsymbol{L}_{ij}\left(\widetilde{\boldsymbol{U}}_{\sigma}^{n}\right),$ (59)
where the operator $\boldsymbol{L}_{ij}$ is defined by (45) with the properly discretized Godunov–Powell source term. The high-order “moments” of the DG solution are updated by evolving the semi-discrete schemes (9) and (12) with the forward Euler method. Note that the scheme (59) ensures $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$, as shown in Theorem 1 later.